OpenStudy (anonymous):

Solve for x: 7x + 15 = 38 twenty-three sevenths fifty-three sevenths 23 53

OpenStudy (texaschic101):

7x + 15 = 38 --- subtract 15 from both sides 7x = 38 - 15 7x = 23 -- divide both sides by 7 x = 23/7

OpenStudy (anonymous):

thankyou

OpenStudy (texaschic101):

sure thing :)